# Data-Centric Foundation Models in Computational Healthcare: A Survey

### v1

@@ 1. Introduction
FM research places a sharp focus on the data-centric perspective (Zha et al., 2023). First, FMs demonstrate the power of scale, where the enlarged model and data size permit FMs to capture vast amounts of information, thus increasing the pressing need of training data quantity (Vaswani et al., 2017). Second, FMs encourage homogenization (Bommasani et al., 2021) as evidenced by their extensive adaptability to downstream tasks. High-quality data for FM training thus becomes critical since it can impact the performance of both pre-trained FM and downstream models. Therefore, addressing key data challenges is progressively recognized as a research priority. In the healthcare system, collecting high-quality records could enable a comprehensive understanding of patient characteristics (imaging, genomics, and lab testing data) (Johnson and Khoshgoftaar, 2023; Singh, 2023; Alfasly et al., 2023). As illustrated, data-centric strategies promise to reshape clinical workflow (Rao et al., 2023; Juluru et al., 2021), enable precise diagnosis (Hunter et al., 2022), and uncover insights into treatment (Chen et al., 2023a).
@@
Medical data challenges have posed persistent obstacles over decades, including multi-modality data fusion (Section 4), limited data volume (Section 5), annotation burden (Section 6), and the critical concern of patient privacy protection (Section 7) (Chen et al., 2021a; Hathaliya and Tanwar, 2020; Huang et al., 2020; Rajpurkar et al., 2022). To respond, the FM era opens up perspectives to advance data-focused AI analytics. Multi-modal FMs, as a concrete example, can offer scalable data fusion strategies for various data formats (Li et al., 2023a; Ding et al., 2023c). Meanwhile, the appealing trait of FM to generate high-quality data can greatly help address data quantity, scarcity, and privacy in the medicine and healthcare community (Tang et al., 2023a; Ding et al., 2023c; Chambon et al., 2022; Lu et al., 2022; Zhao and Bilen, 2023; Van Noorden and Perkel, 2023). To build responsible solutions for healthcare AI, the evolving perspective on AI-human alignment (Gabriel, 2020; Ngo et al., 2022) has become increasingly important. We discuss the necessity of the real-world applications of FMs aligned with human ethics, equity, and societal norms to reduce potential risks in performance assessment, ethical compliance, and patient safety (Liu et al., 2023d; Ozmen Garibay et al., 2023; Liang et al., 2022; Hathaliya and Tanwar, 2020). In the FM era, enabling AI-human alignment further underscores the significance of data focus, motivating us to prioritize the data-centric challenges in the landscape of computational healthcare.
@@
GLIP (Li et al., 2022); CoCa (Yu et al., 2022); FLAVA (Singh et al., 2022); CLAP (Elizalde et al., 2023)] [Medical FMs: ConVIRT (Zhang et al., 2022); BioViL (Boecking et al., 2022);
@@
BiomedJourney (Gu et al., 2023)] ] [Leveraging LLMs [General FMs: Frozen (Tsimpoukelli et al., 2021); SimVLM (Wang et al., 2021);
Flamingo (Alayrac et al., 2022); BLIP-2 (Li et al., 2023c); LLaVA (Liu et al., 2023b);
@@
Visual Med-Alpaca (Shu et al., 2023); Med-Flamingo (Moor et al., 2023b)] ] ] [Quantity
@@
CLIP (Zhang et al., 2023a); GLIP (Yi et al., 2023)] [Medical FM-based: REMEDIS (Azizi et al., 2023);
@@
(Section 6) [Text Annotation [Gilson et al. (Gilson et al., 2022); PathAsst (Sun et al., 2023); DeID-GPT (Liu et al., 2023e);
Med-PaLM 2 (Singhal et al., 2023); Tang et al. (Tang et al., 2023b)] ] [Image Annotation [Zero-shot SAM: SAM (Kirillov et al., 2023); Hu et al. (Hu and Li, 2023);
@@ 2. Foundation Models
As a concrete example of FM, large language models (LLMs) are pre-trained on Internet-scale corpora showing impressive semantic understanding and high-quality text generation (Brown et al., 2020; Touvron et al., 2023a; Chowdhery et al., 2022). In particular, LLMs emphasize the engineering design of input context, also known as the prompt. This design opens the door for human-and-machine interaction, allowing us to feed various prompt queries (e.g., text sequences) into a well-trained FM towards generating desired content or outcomes.
@@ 2.2. Fine-Tuning
The paradigm of pre-training and fine-tuning has dominated deep learning since self-supervised learning enabled large-scale model pre-training. In detail, fine-tuning refers to updating pre-trained model parameters through forward and backward passes and gradient descent on task-specific supervised data. Fine-tuning is a critical technique in constructing medical domain-specific FMs (Lee et al., 2020; Gema et al., 2023; Li et al., 2023e). However, conventional fine-tuning poses significant challenges including high cost, extensive data requirements, and limited generalization to unseen tasks. By contrast, we introduce two essential techniques beyond basic fine-tuning to address those issues: parameter-efficient fine-tuning and instruction tuning.
@@ 2.2.1. Parameter-efficient fine-tuning
Parameter-efficient fine-tuning (PEFT) is a family of fine-tuning approaches that only updates a small set of model parameters while keeping most of the pre-trained weights fixed. These methods alleviate critical challenges caused by updating all parameters of a large FM during full fine-tuning, such as the excessively high cost, overfitting on small downstream datasets, and the risk of catastrophic forgetting (Ding et al., 2023a). PEFT approaches excel in scenarios with limited data availability and effectively retain the valuable knowledge acquired through FM pre-training. We discuss essential PEFT methods. BitFit (Zaken et al., 2021) fine-tunes only the bias parameters of the linear layers in the model. Adapter (Houlsby et al., 2019) inserts lightweight modules into the transformer blocks. Prompt tuning (Lester et al., 2021) prepends trainable tokens to the tokenized input sequence of the first transformer layer. Prefix tuning (Li and Liang, 2021) is like prompt tuning, except that trainable tokens are prepended to the input of all transformer layers. Low-Rank Adaptation (LoRA) (Hu et al., 2021) reparameterizes the weight matrices with multiplications of low-rank matrices. PEFT methods demonstrate potential effectiveness in healthcare (Gema et al., 2023).
@@ 2.3. In-Context Learning
ICL promises to extend powerful LLMs into downstream tasks by injecting contextual information. We introduce essential ICL strategies related to FM applications. Zero-shot is the simplest type of prompt, where we directly assign a task to the model, e.g., “Tell me a joke”. Few-shot gives the model a few demonstrations of the task. For example, “bird = oiseau. cat = chat. dog =”. In this case, we prompt the model to translate the English word “dog” to French by giving two examples. Chain-of-thought (CoT) (Kojima et al., 2022) prompts LLMs to generate the reasoning path (i.e., step-by-step thoughts) for addressing complicated problems. For example, “Q: I had 10 apples. I gave my neighbor 3 apples. Then I bought 4 apples and ate 2. How many apples do I have now? A: Let’s think step by step:”. Based on CoT, self-consistency (Wang et al., 2022b) aggregates a diverse set of reasoning paths to get the most consistent answer. Recently, tree-of-thought (ToT) (Yao et al., 2023) further builds a tree of reasoning paths and applies tree search methods to obtain the best path. Self-refine (Madaan et al., 2023) prompts an LLM to provide feedback for its own outputs and refine itself. We can even prompt an LLM to write prompts.
FMs start to demonstrate potential generalization power in the medical field using ICL techniques (Roy et al., 2023; Lyu et al., 2023). For instance, ChatGPT can effectively translate radiology reports into plain language by using well-designed prompts that inform the model about the structure of the report and contents of each paragraph (Lyu et al., 2023). Also, the general-purpose SAM performs well on abdominal CT organ segmentation when provided point and bounding box prompts in an oracle manner (Roy et al., 2023).
@@ 3.1.2. Adapting via in-context learning (ICL)
ICL has proven to be effective in adapting FMs, especially large language models (LLMs), to a variety of healthcare tasks. With carefully designed task-specific input context (i.e., prompts), the FM can perform well on healthcare tasks without modifying any model parameters. For instance, Nori et al. (Nori et al., 2023) evaluate GPT-4 (OpenAI, 2023) on the United States Medical Licensing Examination (USMLE) without specially crafted prompts. GPT-4 shows its promising zero-shot performance without adding relevant medical context data. Lyu et al. (Lyu et al., 2023) leverage ChatGPT (OpenAI, 2021) to translate radiology reports into plain language for report understanding and translation. The experiments show that by using a clearer and more structured prompt, the overall translation quality can be increased. Roy et al. (Roy et al., 2023) show that SAM (Kirillov et al., 2023) has a great generalization ability in abdominal CT organ segmentation, through its point/bounding box prompting. Deng et al. (Deng et al., 2023) evaluate the zero-shot performance of SAM on tumor segmentation, non-tumor tissue segmentation, and cell nuclei segmentation on whole slide images (WSI), demonstrating that SAM performs well on large connected objects on pathological scans. Chen et al. propose Diagnosis of Thought (DoT) prompting (Chen et al., 2023b) to assist professionals with cognitive distortion detection. DoT diagnoses mental illness by prompting LLMs to sequentially perform subjectivity assessment, contrastive reasoning, and schema analysis.
@@ 3.2.1. Pre-training strategy
The first pre-training strategy is masked language/image modeling, following BERT (Devlin et al., 2018) and masked autoencoder (MAE) (He et al., 2022). For instance, SciBERT (Beltagy et al., 2019) and PubMedBERT (Gu et al., 2021) are pre-trained on multi-domain scientific publications and biomedical domain-specific corpora respectively, based on BERT strategy. BioLinkBERT (Yasunaga et al., 2022) utilizes links between biomedical documents and is pre-trained with masked language modeling and document relation prediction tasks. BioGPT (Luo et al., 2022) is pre-trained on PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/ abstracts following GPT-2 (Radford et al., 2019) for generative language tasks. RETFound (Zhou et al., 2023a) is an FM for retinal image disease detection, pre-trained on a large collection of unannotated retinal images to reconstruct input images with 75% masked patches, following MAE. Similarly, General Expression Transformer (GET) (Fu et al., 2023) is an FM for modeling transcriptional regulation across 213 human cell types. GET is pre-trained to predict the motif binding scores of masked regulatory elements in the input to learn regulatory patterns.
Contrastive learning is another important pre-training strategy for medical FMs. For example, REMEDIS (Azizi et al., 2023) is a medical vision model pre-trained via contrastive learning to extract representative visual features for medical images. Vision-language models such as MedCLIP (Wang et al., 2022c), MI-Zero (Lu et al., 2023b), and PLIP (Huang et al., 2023a), are contrastively pre-trained on domain-specific image-text pairs. They achieve positive performance on zero-shot image classification tasks in radiology and pathology.
@@ 3.2.2. Model initialization
Healthcare FM pre-training benefits from utilizing general FMs as the initial model to leverage their massive information. With proper model initialization, we recognize that much less data and fewer training epochs are needed for domain-specific pre-training of medical FMs. For example, BioBERT (Lee et al., 2020) is pre-trained on PubMed abstracts and PubMed Central (PMC [^2] https://www.ncbi.nlm.nih.gov/pmc/) full-text articles with BERT initialization and demonstrates better performance compared with BERT and previous state-of-the-art models when fine-tuned on three biomedical language tasks. Similarly, BlueBERT (Peng et al., 2019) continues BERT pre-training on PubMed abstracts and MIMIC-III clinical notes (Johnson et al., 2016). Alsentzer et al. (Alsentzer et al., 2019) build clinical text models initialized from BERT and BioBERT, and show a significant improvement on five clinical language tasks. PMC-LLaMA (Wu et al., 2023d) is an open-source language model fine-tuned from LLaMA (Touvron et al., 2023a) on biomedical academic papers. Singhal et al.  (Singhal et al., 2022) apply instruction tuning on Flan-PaLM (Chung et al., 2022) to obtain Med-PaLM model, achieving state-of-the-art performance on a broad range of medical question answering (MedQA) tasks. PubMedCLIP (Eslami et al., 2021), a CLIP model fine-tuned on medical image-text pairs from PubMed articles, shows outstanding performance on medical visual question answering (MedVQA) tasks. However, medical data heterogeneity has long been an obstacle to FM pre-training, generalization, and evaluation. As we see that medical data records reflect the complex nature of human disease complexity in clinics, overcoming this data heterogeneous challenge is crucial to enable FM robustness in real-world healthcare applications (Hathaliya and Tanwar, 2020; Chen et al., 2021a; Rajpurkar et al., 2022; Liang et al., 2022; Huang et al., 2020).

### v2

@@ 1. Introduction
FM research places a sharp focus on the data-centric perspective (Zha et al., 2023). First, FMs demonstrate the power of scale, where the enlarged model and data size permit FMs to capture vast amounts of information, thus increasing the pressing need of training data quantity (Vaswani et al., 2017). Second, FMs encourage homogenization (Bommasani et al., 2021) as evidenced by their extensive adaptability to downstream tasks. High-quality data for FM training thus becomes critical since it can impact the performance of both pre-trained FM and downstream models. Therefore, addressing key data challenges is progressively recognized as a research priority. In the healthcare system, collecting high-quality records could enable a comprehensive understanding of patient characteristics (imaging, genomics, and lab testing data) (Johnson and Khoshgoftaar, 2023; Singh, 2023; Alfasly et al., 2023). As illustrated, data-centric strategies promise to reshape clinical workflow (Rao et al., 2023; Juluru et al., 2021), enable precise diagnosis (Hunter et al., 2022), and uncover insights into treatment (Chen et al., 2023c).
@@
Medical data challenges have posed persistent obstacles over decades, including multi-modality data fusion (Section 4), limited data volume (Section 5), annotation burden (Section 6), and the critical concern of patient privacy protection (Section 7) (Chen et al., 2021a; Hathaliya and Tanwar, 2020; Huang et al., 2020; Rajpurkar et al., 2022). To respond, the FM era opens up perspectives to advance data-focused AI analytics. Multi-modal FMs, as a concrete example, can offer scalable data fusion strategies for various data formats (Li et al., 2023a; Ding et al., 2023c). Meanwhile, the appealing trait of FM to generate high-quality data can greatly help address data quantity, scarcity, and privacy in the medicine and healthcare community (Tang et al., 2023a; Ding et al., 2023c; Chambon et al., 2022; Lu et al., 2022; Zhao and Bilen, 2023; Van Noorden and Perkel, 2023). To build responsible solutions for healthcare AI, the evolving perspective on AI-human alignment (Gabriel, 2020; Ngo et al., 2022) has become increasingly important. We discuss the necessity of the real-world applications of FMs aligned with human ethics, equity, and societal norms to reduce potential risks in performance assessment, ethical compliance, and patient safety (Liu et al., 2023e; Ozmen Garibay et al., 2023; Liang et al., 2022; Hathaliya and Tanwar, 2020). In the FM era, enabling AI-human alignment further underscores the significance of data focus, motivating us to prioritize the data-centric challenges in the landscape of computational healthcare.
@@
GLIP (Li et al., 2022); CoCa (Yu et al., 2022); FLAVA (Singh et al., 2022); CLAP (Elizalde et al., 2023)] [Medical FMs: ConVIRT (Zhang et al., 2022); EchoCLIP (Christensen et al., 2024);
@@
BiomedJourney (Gu et al., 2023); BioViL (Boecking et al., 2022); PRISM (Shaikovski et al., 2024);
MoleculeSTM (Liu et al., 2023c)] ] [Leveraging LLMs [General FMs: BLIP-2 (Li et al., 2023c); SimVLM (Wang et al., 2021);
Flamingo (Alayrac et al., 2022); GPT-4o (OpenAI, 2024); LLaVA (Liu et al., 2023b);
@@
Visual Med-Alpaca (Shu et al., 2023); Med-Flamingo (Moor et al., 2023b);
CheXagent (Chen et al., 2024)] ] ] [Quantity
@@
CLIP (Zhang et al., 2023a); GLIP (Yi et al., 2023); REALM (Zhu et al., 2024); BioRAG (Wang et al., 2024)] [Medical FM-based: REMEDIS (Azizi et al., 2023);
@@
(Section 6) [Text Annotation [Gilson et al. (Gilson et al., 2022); PathAsst (Sun et al., 2023); DeID-GPT (Liu et al., 2023f);
Med-PaLM 2 (Singhal et al., 2023a); MedGraphRAG (Wu et al., 2024);
Tang et al. (Tang et al., 2023b)] ] [Image Annotation [Zero-shot SAM: SAM (Kirillov et al., 2023); Hu et al. (Hu and Li, 2023);
@@ 2. Foundation Models
As a concrete example of FM, large language models (LLMs) are pre-trained on Internet-scale corpora, showing impressive semantic understanding and high-quality text generation (Brown et al., 2020; Touvron et al., 2023a; Bai et al., 2023; OpenAI, 2023). In particular, LLMs emphasize the engineering design of input context, also known as the prompt. This design opens the door for human-machine interaction, allowing us to feed various prompt queries into a well-trained FM towards generating desired content or outcomes.
@@ 2.2. Fine-Tuning
The paradigm of pre-training and fine-tuning has dominated deep learning since self-supervised learning enabled large-scale model pre-training. In detail, fine-tuning refers to updating pre-trained model parameters through forward and backward passes and gradient descent on task-specific supervised data. Fine-tuning is a critical technique in constructing medical domain-specific FMs (Gema et al., 2023; Li et al., 2023e; Saab et al., 2024). However, conventional fine-tuning poses significant challenges including high cost, extensive data requirements, and limited generalization to unseen tasks. By contrast, we introduce two essential techniques beyond basic fine-tuning to address those issues: parameter-efficient fine-tuning and instruction tuning.
@@ 2.2.1. Parameter-efficient fine-tuning
Parameter-efficient fine-tuning (PEFT) is a family of fine-tuning approaches that only updates a small set of model parameters while keeping most of the pre-trained weights fixed. These methods alleviate critical challenges caused by updating all parameters of a large FM during full fine-tuning, such as the excessively high cost, overfitting on small downstream datasets, and the risk of catastrophic forgetting (Ding et al., 2023a). PEFT approaches excel in scenarios with limited data availability and effectively retain the valuable knowledge acquired through FM pre-training. We discuss essential PEFT methods. BitFit (Zaken et al., 2021) fine-tunes only the bias parameters of the linear layers in the model. Adapter (Houlsby et al., 2019) inserts lightweight modules into the transformer blocks. Prompt tuning (Lester et al., 2021) prepends trainable tokens to the tokenized input sequence of the first transformer layer. Prefix tuning (Li and Liang, 2021) is like prompt tuning, except that trainable tokens are prepended to the input of all transformer layers. Low-Rank Adaptation (LoRA) (Hu et al., 2021) reparameterizes the weight matrices with multiplications of low-rank matrices. PEFT methods have demonstrated potential effectiveness in healthcare (Gema et al., 2023).
@@ 2.3. In-Context Learning
ICL promises to extend powerful LLMs into downstream tasks by injecting contextual information. We introduce essential ICL strategies related to FM applications. Zero-shot is the simplest type of prompt, where we directly assign a task to the model. Few-shot gives the model a few demonstrations of the task. For example, “bird = oiseau. cat = chat. dog =”. In this case, we prompt the model to translate the English word “dog” to French by giving two examples. Chain-of-thought (CoT) (Kojima et al., 2022) prompts LLMs to generate the reasoning path (i.e., step-by-step thoughts) for addressing complicated problems. Based on CoT, self-consistency (Wang et al., 2022b) aggregates a diverse set of reasoning paths to get the most consistent answer. Tree-of-thought (ToT) (Yao et al., 2023) further builds a tree of reasoning paths and applies tree search methods to obtain the best path. Self-refine (Madaan et al., 2023) prompts an LLM to provide feedback for its own outputs and refine itself. We can even prompt an LLM to write prompts.
ICL can provide grounded knowledge to compensate for the lag in pre-training data of FMs and the lack of domain-specific knowledge. Retrieval-augmented generation (RAG) is a technique that leverages external knowledge retireval mechanisms to gather additional relevant information for ICL, enhancing the generation quality of FMs (Lewis et al., 2020; Zhao et al., 2024). In RAG, FMs (typically, LLMs) first generate queries based on user input, which are then used to retrieve information from a knowledge base or external documents. The retrieved context is integrated with the input, allowing FMs to produce more informed and accurate outputs.
FMs start to demonstrate generalization power in the medical field using ICL techniques (Roy et al., 2023; Lyu et al., 2023). For instance, ChatGPT can effectively translate radiology reports into plain language by using well-designed prompts that inform the model about the structure of the report and contents of each paragraph (Lyu et al., 2023). Also, the general-purpose SAM performs well on abdominal CT organ segmentation when provided point and bounding box prompts in an oracle manner (Roy et al., 2023). Besides, RAG methods show great potential in enhancing the capabilities of LLMs for medical-related NLP tasks (Zhu et al., 2024; Wu et al., 2024).
@@ 3.1.2. Adapting via in-context learning (ICL)
ICL has proven to be effective in adapting FMs, especially large language models (LLMs), to a variety of healthcare tasks. With carefully designed task-specific input context (i.e., prompts), the FM can perform well on healthcare tasks without modifying any model parameters. For instance, Nori et al. (Nori et al., 2023) evaluate GPT-4 (OpenAI, 2023) on the United States Medical Licensing Examination (USMLE) without specially crafted prompts. GPT-4 shows its promising zero-shot performance without adding relevant medical context data. Lyu et al. (Lyu et al., 2023) leverage ChatGPT (OpenAI, 2021) to translate radiology reports into plain language for report understanding and translation. The experiments show that by using a clearer and more structured prompt, the overall translation quality can be increased. Deng et al. (Deng et al., 2023) evaluate the zero-shot performance of SAM on tumor segmentation, non-tumor tissue segmentation, and cell nuclei segmentation on whole slide images (WSI), demonstrating that SAM performs well on large connected objects on pathological scans. Chen et al. propose Diagnosis of Thought (DoT) prompting (Chen et al., 2023d) to assist professionals with cognitive distortion detection. DoT diagnoses mental illness by prompting LLMs to sequentially perform subjectivity assessment, contrastive reasoning, and schema analysis. Besides, Zhu et al. (Zhu et al., 2024) augment multi-modal electronic health records (EHR) data with a medical knowledge graph via retrieval-augmented generation (RAG).
@@ 3.2.1. Pre-training strategy
The first pre-training strategy is masked language/image modeling, following BERT (Devlin et al., 2018) and masked autoencoder (MAE) (He et al., 2022). For instance, SciBERT (Beltagy et al., 2019) and PubMedBERT (Gu et al., 2021) are pre-trained on multi-domain scientific publications and biomedical domain-specific corpora respectively, based on BERT strategy. BioGPT (Luo et al., 2022) is pre-trained on PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/ abstracts following GPT-2 (Radford et al., 2019) for generative language tasks. RETFound (Zhou et al., 2023a) is an FM for retinal image disease detection, pre-trained on a large collection of unannotated retinal images to reconstruct input images with 75% masked patches, following MAE. Similarly, General Expression Transformer (GET) (Fu et al., 2023) is an FM for modeling transcriptional regulation across 213 human cell types. GET is pre-trained to predict the motif binding scores of masked regulatory elements in the input to learn regulatory patterns.
Contrastive learning is another important pre-training strategy for medical FMs. For example, REMEDIS (Azizi et al., 2023) is a medical vision model pre-trained via contrastive learning to extract representative visual features for medical images. Pai et al. develop an FM for cancer imaging biomarker discovery by contrastively training a convolutional vision encoder (Pai et al., 2024). Vision-language models such as MedCLIP (Wang et al., 2022c), MI-Zero (Lu et al., 2023c), and PLIP (Huang et al., 2023a), are contrastively pre-trained on domain-specific image-text pairs. They achieve positive performance on zero-shot image classification tasks in radiology and pathology.
@@ 3.2.2. Model initialization
Healthcare FM pre-training benefits from utilizing general FMs as the initial model to leverage their massive information. With proper model initialization, we recognize that much less data and fewer training epochs are needed for domain-specific pre-training of medical FMs. For example, BioBERT (Lee et al., 2020) is pre-trained on PubMed abstracts and PubMed Central (PMC [^2] https://www.ncbi.nlm.nih.gov/pmc/) full-text articles with BERT initialization and demonstrates better performance compared with BERT and previous state-of-the-art models when fine-tuned on three biomedical language tasks. PMC-LLaMA (Wu et al., 2023d) is an open-source language model fine-tuned from LLaMA (Touvron et al., 2023a) on biomedical academic papers. Singhal et al.  (Singhal et al., 2022) apply instruction tuning on Flan-PaLM (Chung et al., 2022) to obtain Med-PaLM model, achieving state-of-the-art performance on a broad range of medical question answering (MedQA) tasks. MMedLM (Qiu et al., 2024) is a multi-lingual language model for medicine, further pre-trained from InternLM (Cai et al., 2024) using over 25B medical-related tokens across six languages. PubMedCLIP (Eslami et al., 2021), a CLIP model fine-tuned on medical image-text pairs from PubMed articles, shows outstanding performance on medical visual question answering (MedVQA) tasks. However, medical data heterogeneity has long been an obstacle to FM pre-training, generalization, and evaluation. As we see that medical data records reflect the complex nature of human disease complexity in clinics, overcoming this data heterogeneous challenge is crucial to enable FM robustness in real-world healthcare applications (Hathaliya and Tanwar, 2020; Chen et al., 2021a; Rajpurkar et al., 2022; Liang et al., 2022; Huang et al., 2020).